OpenStudy (anonymous):

Let X be a discrete random variables with PMF given by : P_X(x)= x/15 ,{x=1,2,3,4,5} 0 , otherwise b)let Y=(X-3)^2, find range Y, S_Y , PMF of Y

OpenStudy (bahrom7893):

@amistre64

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

P_X(x)= x/15 ,{x=1,2,3,4,5} 0 , otherwise so to find range using Y y= (X-3)^2 ((x/15)-3)^2 y= (1/15 -3)^2 (-44/45)^2 to 0

OpenStudy (anonymous):

am I right?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

x takes the values 1,2,3,4,5 y=(x-3)^2 plug those x values into the above equation

OpenStudy (zarkon):

for the range(support)

OpenStudy (anonymous):

oh, it is X, not x

OpenStudy (zarkon):

yes...so X takes the values 1,2,3,4,5 Y=(X-3)^2 plug those X values into the above equation :)

OpenStudy (zarkon):

ok...do what I wrote

OpenStudy (anonymous):

(1-3)^2,(2-3)^2,(3-3)^2,(4-3)^2,(5-3)^3 4,1,0,1,4

OpenStudy (zarkon):

so 0,1,4

OpenStudy (zarkon):

that is the range (or support)

OpenStudy (zarkon):

now you can find the pmf

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so y=(x-3)^2 do we solve for x?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[P_{Y}(0)=P_{X}(3)\] \[P_{Y}(1)=P_{X}(2)+P_{X}(4)\] \[P_{Y}(4)=P_{X}(1)+P_{X}(5)\]

OpenStudy (anonymous):

can you explain me how you got that?

OpenStudy (zarkon):

\[P_{Y}(1)=P(Y=1)=P(X=2\text{ or }X=4)\] \[=P(X=2)+P(X=4)=P_{X}(2)+P_{X}(4)\]

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

still figuring out the last post

OpenStudy (zarkon):

ok

OpenStudy (anonymous):

I understand you sum because it is 'or' but why is P(y=1) = P(x=2 or x=4)

OpenStudy (zarkon):

\[Y=(X-3)^2\] \[1=(2-3)^2\] \[1=(4-3)^2\] if I tell you Y=1 then either X=2 or X=4

OpenStudy (anonymous):

got it

OpenStudy (zarkon):

good

OpenStudy (anonymous):

so that's just P(Y=1) do we find Y=1,2,3,4,5?

OpenStudy (zarkon):

Y only takes the values 0,1,4

OpenStudy (anonymous):

got it , thanks; I wish you were my probability professor